# SoK: Exploring the Potential of Large Language Models for Improving Digital Forensic Investigation Efficiency

### v1

@@ 4.5. Autonomous Agents
The evolution of LLMs, with their capacity to generate information and communicate in a manner resembling human interaction, has led to the development of autonomous agents. The expectation is that these agents will effectively execute a wide array of tasks, capitalising on the human-like capabilities inherent in LLMs (Wang et al., 2023d). These autonomous agents follow a four-stage architecture, encompassing profiling, memory, planning, and action. Profiling defines the agent’s role, privileges, domain, and expertise (Wang et al., 2023d). Memory stores information about tasks and profile data relevant to the environment. Planning involves breaking down given tasks into subtasks and solving them individually. The action stage is the final phase where all decisions and subtasks are translated into actions executed by the agent. Zhang et al. (2023a) developed a framework designed to facilitate collaboration between AGI agents and humans. This framework enables planning and communication for specific tasks, leveraging the capabilities of LLMs Similarly, AgentSims, ToolBench, GameGPT, ChatDev, Voyager, and RecMind represent a diverse array of autonomous AGI agents developed with distinct goals and objectives (Wang et al., 2023b, e; Lin et al., 2023; Qin et al., 2024; Chen et al., 2023b; Qian et al., 2023). Certainly, AutoGen stands out as a multiagent framework with the capability to autonomously perform tasks or collaborate with human feedback. This flexibility makes it a versatile tool for various applications (Wu et al., 2023a).
4.6. Limitations and Risks
@@
As explored in the preceding sections, it appears that LLMs possess a vast range of capabilities. However, it is crucial to acknowledge that they are not without limitations and risks. In multimodal LLMs, it is a common issue that they are over-reliant (Wolf et al., 2024). Hadi et al. (2023) highlighted significant drawbacks associated with LLMs, including issues such as bias, explainability challenges, reasoning errors, logical errors, hallucinations, vulnerability to prompt injections, and spelling and grammar errors. These limitations underscore the importance of a cautious and critical approach when utilising LLMs in various applications. Additionally, the literature shows limitations in LLMs, including statistical inconsistency, the absence of emotional attributes in linguistic responses, and challenges related to fact verification (Tang et al., 2023; Fröhling and Zubiaga, 2021). These factors contribute to a comprehensive understanding of the constraints and potential shortcomings when working with LLMs. Thapa et al. (2023) contend that while LLMs can indeed diminish the time and costs associated with annotation tasks, they fall short of completely supplanting human annotation. This is because they struggle with intricate linguistic constructions, such as idioms, irony, sarcasm, and metaphor, which can potentially impact the precision of annotations.
@@
Similar limitations are associated with MLLMs. Issues such as over-reliance on training data, sensitivity to word order in the input prompts, and vulnerability to prompts containing extra knowledge represent challenges in the capabilities of MLLMs (Qi et al., 2023). There are several more concerns associated with LLMs including restricted input and output text lengths, limited comprehension of syntax, ethical considerations with the generated information, constraints with multilingual capabilities, elevated costs associated with training and maintenance, inadequate understanding of human behaviours, and limited ability to learn incrementally (Hadi et al., 2023; Chiang and yi Lee, 2023; Yang et al., 2023).
@@
Despite their considerable capabilities, LLMs are not without risks. Bommasani et al. (2022); Lund and Wang (2023); Rahman and Santacana (2023) provide comprehensive overviews of risks linked to LLMs. These include the homogenisation of outcomes, wherein defects or biases from the foundation model are inherited by all downstream models. There is also the risk of a monopoly for the owners of the foundation models, leading to increased centralisation of power in a single entity. Ethical and legal concerns are intertwined with privacy and intellectual property issues. Additionally, there are economic and environmental impacts, raising concerns about the potential displacement of human workers. Furthermore, inequity and misuse of LLMs, such as the creation of deepfakes and their application in criminal and unethical activities, pose additional challenges. Given that LLMs do not inherently prioritise the accuracy of information, Bender et al. (2021) have underscored the risk of generating social turbulence, especially, when employed on social media platforms. Additionally, the use of LLMs is associated with significant costs, leading to a direct environmental impact due to their substantial energy consumption (Rillig et al., 2023).
@@
The risks associated with LLMs are predominantly emphasised within Information Communication and Technology (ICT) and cyberspace. Primary concerns include the disclosure of personal information, the generation of malicious text, and the creation of malicious code (Derner and Batistič, 2023).
@@
In response to the increasing prominence of LLMs and Artificial General Intelligence (AGI), the EU Artificial Intelligence Act (AIA) presently classifies AI systems into four risk categories based on broad fields of application. However, this categorisation approach may result in inaccurate risk estimation and ineffective enforcement. To rectify this issue, the authors advocate for a risk assessment model that aligns the risk categories with specific AI scenarios. This model integrates the AIA with the risk assessment approach employed by the Intergovernmental Panel on Climate Change (IPCC) and relevant literature, facilitating a more precise estimation of the magnitude of AI-related risks (Novelli et al., 2023).
@@
The Beyond the Imitation Game benchmark (BIG-bench), serves as an evaluation framework for LLMs. It encompasses 204 distinct language-related tasks, spanning contextual and context-free question-answering, reading comprehension, etc. (Srivastava et al., 2023). It is acknowledged that the challenge of social biases and English language dependency persists across nearly all LLMs.
5. Large Language Models For Digital Forensics
@@
Section 5 summarises existing work with LLMs in DF, the feasibility of employing them, and potential future directions. As discussed in Sections 3 and 4, despite the widespread use of LLMs in various fields to enhance the efficiency and accuracy of tasks within specific domains, their application in the field of DF is still relatively new.
@@
Conducting a thorough analysis of the utilisation of LLMs in conjunction with the stages of the DF process model, as highlighted in Section 2.1, proves to be a valuable undertaking.
5.1. Incident Recognition Phase
@@
In the initial phase of Casey’s DF process model, which delineates the recognition of an incident, LLMs can serve as a valuable detection mechanism. In cases of cybercrime, the primary artefacts often involve data logs, data dumps and network dumps. Fine-tuning an LLM to monitor text-based logs and related files enables it to discern and identify potential or ongoing incidents within the environment. In network-related activities, anomaly detection plays a pivotal role in initiating an incident response. Various existing anomaly detection techniques are employed in systems for this purpose. Leveraging its capability to identify patterns in a series of text data sets, LLMs exhibit potential as an Intrusion Detection System (IDS) within such systems (Li et al., 2023c).
5.2. Collection Phase
@@
5.3. Preservation/Acquisition Phase
Preserving evidence is centred on maintaining integrity. To achieve this, various tools such as EnCase and FTK Imager have been employed, aiding investigators in streamlining their work processes (Shah et al., 2017). In the context of preserving disk evidence, it becomes feasible for an investigator to articulate their requirements in natural language, LLM to generate source code tailored to the specific needs. Indeed, LLMs specialised in code generation, such as StarCoder, Code LLaMA, and others, can be fine-tuned and retrained for specific tasks, including those related to preserving disk evidence through customised code generation. In certain instances, gathering live data for forensic investigations becomes crucial, particularly data collected at the crime scene. For this purpose, investigators may leverage DFaaS platforms like Hansken. Hansken possesses the capability to amalgamate custom extraction APIs for data extractions, and these APIs can be developed using code-generative LLMs (van Beek et al., 2015). This approach enhances the adaptability and efficiency of the investigative process.
As stated in Section 4.5, the automation of code generation and unit testing can be facilitated by autonomous agents that utilise LLMs as their core. AutoGen, being an open-source framework, provides the means to develop AI agents tailored for specific tasks. These AutoGen agents are not only customizable and conversational but can also operate in diverse modes, employing combinations of LLMs, human inputs, and various tools (Wu et al., 2023a). Positively, automated agents, particularly those developed within frameworks like AutoGen, can be effectively employed in the preservation phase. Their capabilities will enable the automated preservation within this phase, streamlining and enhancing efficiency in the preservation of digital evidence (Wu et al., 2023a).
5.4. Examination Phase
@@
5.5. Analysis Phase
@@
The analysis phase involves comprehending the incident and arriving at a conclusive understanding based on the information gathered during the examination phase. As also highlighted in Section 2.3, it has been demonstrated that LLMs are effective in case analysis (Henseler and van Beek, 2023). The use of MLLMs, which possess the capability to interpret images, broadens the scope for analysing a crime case more comprehensively. LLMs can be specifically fine-tuned for the analysis of various data types, including log files, email contents, chat transcripts, call records, file metadata, hex dumps, memory dumps, and registry hives. Incorporating contents such as event logs, timestamps, and network traffic captures further enables the effective recreation of incidents by correlating each data set with the assistance of LLMs. Also, audio and video-specific MLLMs can assist in analysing the contents within these formats. This specialised capability has the potential to significantly save time for investigators in their analyses of audio and video data during investigations.
@@
The use of automated agents can distribute the analysis workload. Simultaneously, leveraging Augmented Large Language Models (ALLM) and Retrieval Augmented Generation (RAG) techniques can enhance real-time knowledge searching, thereby improving the accuracy of analysis and decision-making processes  (Lewis et al., 2020).
5.6. Reporting Phase
@@
The quality and validity of evidence, along with the thoroughness of the analysis, are encapsulated in the final report. The reporting phase holds significant weight, as the entire judgement may hinge on this crucial stage. Notably, DF is experiencing heightened scrutiny with regard to the quality of the reports, emphasising the importance of precision and clarity in this phase (Karie et al., 2019). As pointed out by Champod et al. (2016), there is no standard framework for evaluating and reporting scientific findings to authorities and stakeholders. To provide assistance and alleviate scrutiny, incorporating LLMs for report creation is a viable solution. While by definition, LLMs do not produce deterministic outputs, they could be retrained using standards such as ISO/IEC 27043:2015, the international standard for the DF investigation process (Valjarević et al., 2016). A first feasibility study has already been released by Michelet and Breitinger (2023). Naturally, LLMs could also help to automate forensic reports and alternative formats (i.e., not MS Word) such as HTML or Latex.
5.7. Other Possibilities
@@
In general, LLMs can be effectively utilised in each phase of the DF process model. By integrating the DF process life cycle with automated agents, it becomes feasible to establish a fully automated DF process throughout an investigation. For instance, if automated agents are assigned specific roles and maintain synchronisation within each agent, employing frameworks such as AutoGen, there is a substantial potential to expedite the completion of one or multiple DF cases compared to relying solely on human agents. This approach enhances efficiency, speed, and consistency in the DF investigative process.
@@
Scanlon et al. (2023a) highlights that LLMs can play a significant role in DF teaching scenarios. This involvement extends to activities such as storyboarding, synthetic content creation, and synthetic character profiling. Additionally, fine-tuned LLMs prove to be valuable tools for establishing evaluation criteria for DF agents, facilitating the creation of exams, measuring the accuracy and efficiency of investigations through an automated weighing system, and effectively presenting a case to non-technical or non-domain individuals, such as judges.
5.8. Challenges for LLMs in Digital Forensics

### v2

@@ 3.5 Autonomous Agents
The evolution of LLMs, with their capacity to generate information and communicate in a manner resembling human interaction, has led to the development of autonomous agents. The expectation is that these agents will effectively execute a wide array of tasks, capitalising on the human-like capabilities inherent in LLMs [119]. These autonomous agents follow a four-stage architecture, encompassing profiling, memory, planning, and action. Profiling defines the agent’s role, privileges, domain, and expertise [119]. Memory stores information about tasks and profile data relevant to the environment. Planning involves breaking down given tasks into subtasks and solving them individually. The action stage is the final phase where all decisions and subtasks are translated into actions executed by the agent. Zhang et al. [120] developed a framework designed to facilitate collaboration between AGI agents and humans. This framework enables planning and communication for specific tasks, leveraging the capabilities of LLMs Similarly, AgentSims, ToolBench, GameGPT, ChatDev, Voyager, and RecMind represent a diverse array of autonomous AGI agents developed with distinct goals and objectives [121, 122, 123, 124, 125, 126]. Certainly, AutoGen stands out as a multiagent framework with the capability to autonomously perform tasks or collaborate with human feedback. This flexibility makes it a versatile tool for various applications [127].
In addition to the AutoGen framework, AgentLite, Camel and CrewAI are each similar LLM-based agent framework architectures. These platforms are distinguished by their support for task decomposition, multi-agent orchestration, and adaptable reasoning. Notably, AgentLite and CrewAI facilitate work delegation functionalities, augmenting their utility in various operational contexts [128, 129].
3.6 Retrieval-augmented Generation
The content generated by LLMs is highly dependent on the extensive text-based datasets on which they are trained. These datasets may contain vast amounts of information utilised by the LLMs. However, maintaining up-to-date knowledge within LLMs is challenging, as fine-tuning or retraining a model is often extremely costly and resource-intensive. Retrieval-augmented Generation (RAG) is a technique designed to address this LLM knowledge gap by retrieving information from external sources and integrating it with the model’s internal representations [130].
A major advantage of RAG is its ability to reduce the hallucination problem in LLMs, enabling them to generate more accurate and current information [131, 132]. The architecture of RAG consists of a knowledge base and a retriever model. The retriever model converts input prompts and content from the knowledge base into vectors. The user prompts are then appended with the most relevant content from the knowledge base, and this augmented prompt is sent to the base model to generate a more accurate response [132].
3.7 Limitations and Risks
@@
As explored in the preceding sections, LLMs appear to possess a vast range of capabilities. However, it is crucial to acknowledge that they are not without limitations and risks. In multimodal LLMs, it is a common issue that they are over-reliant [133]. Hadi et al. [134] highlighted significant drawbacks associated with LLMs, including issues such as bias, explainability challenges, reasoning errors, logical errors, hallucinations, vulnerability to prompt injections, and spelling and grammar errors. These limitations underscore the importance of a cautious and critical approach when utilising LLMs in various applications. Additionally, the literature shows limitations in LLMs, including statistical inconsistency, the absence of emotional attributes in linguistic responses, and challenges related to fact verification [135, 136]. These factors contribute to a comprehensive understanding of the constraints and potential shortcomings when working with LLMs. Thapa et al. [137] contend that while LLMs can indeed diminish the time and costs associated with annotation tasks, they fall short of completely supplanting human annotation. This is because they struggle with intricate linguistic constructions, such as idioms, irony, sarcasm, and metaphor, which can potentially impact the precision of annotations.
@@
Similar limitations are associated with MLLMs, such as over-reliance on training data, sensitivity to word order in the input prompts, and vulnerability to prompts containing extra knowledge [138]. There are several more concerns associated with LLMs including restricted input and output text lengths, limited comprehension of syntax, ethical considerations with the generated information, constraints with multilingual capabilities, elevated costs associated with training and maintenance, inadequate understanding of human behaviours, and limited ability to learn incrementally [134, 139, 140].
@@
Despite their considerable capabilities, LLMs are not without risks. Bommasani et al. [141], Lund and Wang [142], Rahman and Santacana [143] provide comprehensive overviews of risks linked to LLMs. These include the homogenisation of outcomes, wherein defects or biases from the foundation model are inherited by all downstream models. There is also the risk of a monopoly for the owners of the foundation models, leading to increased centralisation of power in a single entity. Ethical and legal concerns are intertwined with privacy and intellectual property issues. Additionally, there are economic and environmental impacts, raising concerns about the potential displacement of human workers. Furthermore, inequity and misuse of LLMs, such as the creation of deepfakes and their application in criminal and unethical activities, pose additional challenges. Given that LLMs do not inherently prioritise the accuracy of information, Bender et al. [144] have underscored the risk of generating social turbulence, especially, when employed on social media platforms. Additionally, the use of LLMs is associated with significant costs, leading to a direct environmental impact due to their substantial energy consumption [145].
@@
The risks associated with LLMs are predominantly emphasised within Information Communication and Technology (ICT) and cyberspace. Primary concerns include the disclosure of personal information, the generation of malicious text, and the creation of malicious code [146].
@@
In response to the increasing prominence of LLMs and Artificial General Intelligence (AGI), the EU Artificial Intelligence Act (AIA) presently classifies AI systems into four risk categories based on broad fields of application. However, this categorisation approach may result in inaccurate risk estimation and ineffective enforcement. To rectify this issue, the authors advocate for a risk assessment model that aligns the risk categories with specific AI scenarios. This model integrates the AIA with the risk assessment approach employed by the Intergovernmental Panel on Climate Change (IPCC) and relevant literature, facilitating a more precise estimation of the magnitude of AI-related risks [147].
@@
The Beyond the Imitation Game benchmark (BIG-bench), serves as an evaluation framework for LLMs. It encompasses 204 distinct language-related tasks. These span contextual and context-free question-answering, reading comprehension, logical reasoning, etc. [148]. It is acknowledged that the challenge of social biases and English language dependency persists across nearly all LLMs.
When employing LLM-based agents, it is imperative to address challenges associated with LLM-based multi-agent frameworks as well. Handling many defined agents may necessitate substantial computational resources and memory, thereby mandating high-end computing infrastructure for seamless operations. Furthermore, the absence of a standardised comprehensive benchmarking system to evaluate the behaviour of such agents underscores the limitations inherent in the development of LLM-based multi-agent systems. These challenges underscore the need for further research and refinement in this domain to enhance the efficiency and effectiveness of LLM-based multi-agent frameworks [149].
Similar to these risks, the Open Web Application Security Project (OWASP) has identified ten major risk factors related to LLMs. These risks include prompt injection, insecure output handling, training data poisoning, model denial of service, supply chain vulnerabilities, sensitive information disclosure, insecure plugins, excessive agency, overreliance, and model theft. Despite these threat factors, OWASP also emphasised the need for regulatory bodies to oversee LLMs across various domains and recommended the implementation of risk management programs that incorporate the checklist provided by OWASP [^2] https://owasp.org/www-project-top-10-for-large-language-model-applications.
4 Large Language Models For Digital Forensics
@@
Section 4 summarises existing work with LLMs in DF, the feasibility of employing them, and potential future directions. As discussed in Sections 2 and 3, despite the widespread use of LLMs in various fields to enhance the efficiency and accuracy of tasks within specific domains, their application in the field of DF is still relatively new.
@@
Conducting a thorough analysis of the utilisation of LLMs in conjunction with the stages of the DF process model, as highlighted in Section 1.1.1, proves to be a valuable undertaking.
4.1 Incident Recognition Phase
@@
In the initial phase of Casey’s DF process model, which delineates the recognition of an incident, LLMs can serve as a valuable detection mechanism [150]. In cybercrime cases, the primary artefacts often involve data logs, data dumps and network dumps. Fine-tuning an LLM to monitor text-based logs and related files enables it to discern and identify potential or ongoing incidents within the environment. In network-related activities, anomaly detection plays a pivotal role in initiating an incident response. Various existing anomaly detection techniques are employed in systems for this purpose. Leveraging its capability to identify patterns in a series of text data sets, LLMs exhibit potential as an Intrusion Detection System (IDS) within such systems [151].
4.2 Collection Phase
@@
4.3 Preservation/Acquisition Phase
Preserving evidence is centred on maintaining integrity. To achieve this, various tools such as EnCase and FTK Imager have been employed, aiding investigators in streamlining their work processes [152]. In the context of preserving disk evidence, it becomes feasible for an investigator to articulate their requirements in natural language, LLM to generate source code tailored to the specific needs. Indeed, LLMs specialised in code generation, such as StarCoder, Code LLaMA, and others, can be fine-tuned and retrained for specific tasks, including those related to preserving disk evidence through customised code generation. In certain instances, gathering live data for forensic investigations becomes crucial, particularly data collected at the crime scene. For this purpose, investigators may leverage DFaaS platforms like Hansken. Hansken possesses the capability to amalgamate custom extraction APIs for data extractions, and these APIs can be developed using code-generative LLMs [29]. This approach enhances the adaptability and efficiency of the investigative process.
As stated in Section 3.5, the automation of code generation and unit testing can be facilitated by autonomous agents that utilise LLMs as their core. AutoGen, being an open-source framework, provides the means to develop AI agents tailored for specific tasks. These AutoGen agents are not only customizable and conversational but can also operate in diverse modes, employing combinations of LLMs, human inputs, and various tools [127]. Positively, automated agents, particularly those developed within frameworks like AutoGen, can be effectively employed in the preservation phase. Their capabilities will enable the automated preservation within this phase, streamlining and enhancing efficiency in the preservation of digital evidence [127].
4.4 Examination Phase
@@
The use of LAMs and VoT techniques in the examination phase can significantly enhance the efficiency of an investigation. Since LAMs and VoT specifically focus on task manipulation, investigators can offload some examination work to an LAM, which will then generate the final results from a series of subtasks. This approach can allow investigators to focus on higher-level analysis and decision-making, thereby streamlining the overall investigative process.
4.5 Analysis Phase
@@
The analysis phase involves comprehending the incident and arriving at a conclusive understanding based on the information gathered during the examination phase. As also highlighted in Section 1.1.3, it has been demonstrated that LLMs are effective in case analysis [30]. The use of MLLMs, which possess the capability to interpret images, broadens the scope for analysing a crime case more comprehensively. LLMs can be specifically fine-tuned for the analysis of various data types, including log files, email contents, chat transcripts, call records, file metadata, hex dumps, memory dumps, and registry hives. Incorporating contents such as event logs, timestamps, and network traffic captures further enables the effective recreation of incidents by correlating each data set with the assistance of LLMs. Also, audio and video-specific MLLMs can assist in analysing the contents within these formats. This specialised capability can significantly reduce the time investigators spend analysing audio and video data during investigations.
@@
The use of automated agents can distribute the analysis workload effectively. Simultaneously, leveraging Augmented Large Language Models (ALLMs) and RAG techniques can enhance real-time knowledge retrieval, thereby improving the accuracy of analysis and decision-making processes [153]. For example, integrating a source of intelligence with an RAG system can assist investigators in connecting the dots during a DF investigation.
4.6 Reporting Phase
@@
The quality and validity of evidence, along with the thoroughness of the analysis, are encapsulated in the final report. The reporting phase holds significant weight, as the entire judgement may hinge on this crucial stage. Notably, DF is experiencing heightened scrutiny about the quality of the reports, emphasising the importance of precision and clarity in this phase [154]. As pointed out by Champod et al. [155], there is no standard framework for evaluating and reporting scientific findings to authorities and stakeholders. To provide assistance and alleviate scrutiny, incorporating LLMs for report creation is a viable solution. While by definition, LLMs do not produce deterministic outputs, they could be retrained using standards such as ISO/IEC 27043:2015, the international standard for the DF investigation process [156]. A first feasibility study has already been released by Michelet and Breitinger [157]. Naturally, LLMs could also help to automate forensic reports and alternative formats (i.e., not MS Word) such as HTML or Latex.
4.7 Other Possibilities
@@
In general, LLMs can be effectively used in each phase of the DF process model. By integrating the life cycle of the DF process with automated agents, it becomes feasible to establish a fully automated DF process throughout an investigation. For instance, if automated agents are assigned specific roles and maintain synchronisation within each agent, employing frameworks such as AutoGen, there is a substantial potential to expedite the completion of one or multiple DF cases compared to relying solely on human agents. This approach improves efficiency, speed, and consistency in the DF investigation process.
@@
Scanlon et al. [13] highlights that LLMs can play an important role in DF teaching scenarios. This involvement extends to activities such as storyboarding, synthetic content creation, and synthetic character profiling. Additionally, fine-tuned LLMs prove to be valuable tools for establishing evaluation criteria for DF agents, facilitating the creation of exams, measuring the accuracy and efficiency of investigations through an automated weighing system, and effectively presenting a case to non-technical or non-domain individuals, such as judges.
Additionally, to provide a comprehensive understanding of the potential utilisation of LLMs in DF, Table LABEL:DF_example_prompts elucidates sample DF functionalities within the framework of the National Institute of Standards and Technology (NIST) Computer Forensics Tool Testing Program (CFTT), highlighting the usability of LLMs and example prompts. The CFTT project establishes overarching specifications to assess the capabilities of DF tools, a framework adopted by numerous prominent free and commercial DF tools  [^3] https://www.nist.gov/itl/ssd/software-quality-group/computer-forensics-tool-testing-program-cftt.
Table 3: DF functionalities by CFTT highlighting the usability of LLMs and example prompts
CFTT Functionality DF Phase(s) Usable LLMs/Agent Frameworks Example Prompt
Cloud Data Extraction Acquisition LLaMA (Fine-Tuned), Code Llama, StarCoder, AutoGen or CrewAI Retrieve all the data inside given S3 bucket by using given credentials
Deleted File Recovery Specs Acquisition, Examination LLaMA (Fine-Tuned), Code Llama, StarCoder, AutoGen or CrewAI Find all the deleted files from the X disk image and recover them to Y location
Disk Imaging Acquisition LLaMA (Fine-Tuned), Code Llama, StarCoder, AutoGen or CrewAI Get a full disk image from this computer and save it in Z location
Forensic File Carving Acquisition, Examination LLaMA (Fine-Tuned), Code Llama, StarCoder, AutoGen or CrewAI Find all the deleted PDF files from X disk image
Forensic Media Preparation Acquisition LLaMA (Fine-Tuned), Code Llama, StarCoder, AutoGen or CrewAI Prepare the given X device for new investigation
String Search Examination LLaMA (Fine-Tuned), Code Llama, StarCoder, AutoGen or CrewAI Search all the files containing the email of mail@test.com
Mobile Devices Examination LLaMA (Fine-Tuned), LLaVa (Fine-Tuned), Code Llama, StarCoder, AutoGen or CrewAI Find all the photos taken with a computer within the last 3 months
MS Windows Registry Examination LLaMA (Fine-Tuned), Code Llama, StarCoder, AutoGen or CrewAI Find information about the users from a given Windows disk image
SQLite Databases Examination LLaMA (Fine-Tuned), Code Llama, StarCoder, AutoGen or CrewAI Find the access time of user Y to application X using given SQLite databases
Also, to explore the applicability of LLMs to DF, a simple matrix was defined in Table LABEL:potential_LLMsin_DFphases. This matrix showcases the high, medium, and low potential DF phases where LLMs can be utilised. The usability rankings are based on factors such as the extent of human intervention and effort required, the prospect of automation, and the applicability of human expert knowledge.
{longtblr}
[ caption = Potential of LLMs in Digital Forensic Phases, label = potential_LLMsin_DFphases, ] width = colspec = Q[130]Q[490]Q[170], cells = c, hlines, vlines, Potential & Description DF Phases
High Demands minimal human expert knowledge. Exclusively deals with digital evidence. Fully amenable to automation. Examination, Analysis, Reporting
Medium Necessitates human expert knowledge, though it is not a critical component. Focuses on handling digital evidence, with minimal involvement of physical evidence. Automation is feasible and can be implemented. Preservation
Low Requires a high level of human expert knowledge. Often requires identifying/handling physical evidence. Minimal or no scope for automation. Collection
5 Challenges and Risks
This section discusses the challenges and risks of using LLMs in DF. Despite their promising potential, there are significant risk factors to consider. These risks can have severe consequences for DF if not adequately identified and considered in the DF process.
5.1 Challenges for LLMs in Digital Forensics